Solve the Following problems by using the subtitution method.
1.The units digit of a two-digit number is 3 more than its tens digit. If the number is divided by the sum of its digit's, the quotient is 4 and the remainder is 3. Find the number.

2.Paul's investment is 10 000 less than thrice his sister's investment. If their total inestment amounts to 150 000, how much is Paul's investent?

3.Working together, Alleli and gemma can finnish their math project in 5 hours. Alleli can finish it alone twice as much time as gemma can do the job. How long will it take gemme to finnish the project alone?

PLSSS I NEED YOUR HELP, BECAUSE I DONT KNOW HOW TO DO THIS THANK YOU FOR HELPING ME TO SOLVE THIS MY ASSIGNMENT. THANK YOU :)) PLSSS I NEED THE ANSWER NOW AND PLSS INCLUDE THE SOLUTION BY EACH NUMBER THAN YOU

Answer by mananth(16949)

1.The units digit of a two-digit number is 3 more than its tens digit. If the number is divided by the sum of its digit's, the quotient is 4 and the remainder is 3. Find the number.
y-unit digit
x ten's digit
y=x+3........1
(10x+y+3/(x+y)) =4
10x+y+3=4(x+y)
10x+y+3=4x+4y
6x-3y=-3
/3
2x-y=-1
substitute y
2x-x-3=-1
x=2
plug x in (1)
y=2+3
y=5
Number = 25


.Paul's investment is 10 000 less than thrice his sister's investment. If their total inestment amounts to 150 000, how much is Paul's investent?
Sister = x
Paul = y
y=3x-10000......1
x+y= 150000......2
substitute y in (2)
3x-10000+x=150000
4x=160000
x=$40000---- sister's investment
Paul's investment = 3x-10000===> 120000-10000=$ 110,000


3.Working together, Alleli and gemma can finnish their math project in 5 hours. Alleli can finish it alone twice as much time as gemma can do the job. How long will it take gemme to finnish the project alone?
Gemma takes x hours
so she does 1/x of job in 1 hour
Alleli takes 2x hours
so she does 1/2x of the job in 1 hour
Together they do 1/5 of the job in 1 hour
1/x + 1/2x= 1/5
LCD = 10x multiply equation by the LCD
10+5=2x
2x=15
x= 7.5 hours time taken by Gemma alone
m.ananth@hotmail.ca